If the area of the yard is 256 ft.2 and the walkway is 4 ft. wide, what is the area of the walkway?

Found 2 solutions by solver91311, stanbon:
Answer by solver91311(24713)



That would depend on exactly where the walkway is in relationship to the yard. It would depend on whether the 256 square foot area includes the walkway or not. It would depend on the shape of the yard -- is it a square shaped yard (16 feet by 16 feet) or some other shape? If you are depending on someone having a diagram for this problem, good luck. If you think this is the Psychic Hot Line, step back and take another think.

John


Answer by stanbon(75887)

If the area of the yard is 256 ft.2 and the walkway is 4 ft. wide, what is the area of the walkway?
----------------------
Draw the picture.
You have a rectangle inside a larger rectangle.
---
Equations:
The smaller rectangle has area xy = 256 sq ft.
The larger rectangle has area (x+8)(y+8)
= xy +8x + 8y + 64
----
Area of the walkway = larger area - smaller area
---
= xy+8x+8y+64-256
---
Substitute for xy:
= 256 + 8(x+y) - 256
---
= 8(x+y) sq. ft.
==========================
Cheers,
Stan H.
==========================